OpenStudy (anonymous):

Show that if the curve y=f(x) has a maximum stationary point at x=a, then the curve y=1/f(x) has a minimum stationary point at x=a as long as f(a) does not equal 0. Stationary point is when the derivative of the function=0.

OpenStudy (mathmale):

"Stationary point" is not a commonly seen vocabulary word in calculus, so you might want to (1) look it up yourself and (2) share its definition with the rest of us.

OpenStudy (anonymous):

Stationary point is when the derivative of the function=0. Sorry for not clarifying

OpenStudy (irishboy123):

@chau88 does have a point

OpenStudy (anonymous):

@SolomonZelman, Thank you very much for your answer. I believe that I missed out the numerator f'(x) when I was differentiating y. However, I don't really understand why 1/max=min. What I tried to do was second differentiate it and show that f''(x)>0 and therefore the function is max. But I realised that it was too hard to differentiate \[y'=f'(x)/f(x)^{2}\] Please advise. Thank!

OpenStudy (loser66):

Let \(g(x) = \dfrac{1}{f(x)}\) We need show 1) a is a critical point of g(x) 2) g(a) is minimum. 1) a is a critical point of g(x) : At x =a, \(f'(a) =0\) and we know that f(a) max at a. Now, \(g'(x) = \dfrac{-f'(x)}{f^2(x)}\implies g'(a) = \dfrac{-f'(a)}{f^2(a)}\) We have \(f(a) \neq 0\), hence g'(a) defined Moreover, f'(a) =0, Hence g'(a) =0. That shows a is a critical point of g(x). 2) g(a) is minimum of g(x) Look at the chart for f(x)